OpenStudy (anonymous):

A tractor was purchased for $30,000 six years ago. What is the value of the tractor today if the depreciation rate was 2.5%?

jimthompson5910 (jim_thompson5910):

Hint: FV = PV*(1-r)^t FV = 30000*(1-0.025)^6 FV = ??